# Compressed and Smooth Latent Space for Text Diffusion Modeling

### v1

@@ Abstract
Autoregressive language models dominate modern text generation, yet their sequential nature introduces fundamental limitations: decoding is slow, and maintaining global coherence remains challenging. Diffusion models offer a promising alternative by enabling parallel generation and flexible control; however, their application to text generation is hindered by the high dimensionality of token-level representations. We introduce Cosmos, a novel approach to text generation that operates entirely in a compressed, smooth latent space tailored specifically for diffusion. This space is learned using an autoencoder trained simultaneously for token-level reconstruction and alignment with frozen activations from a pretrained language encoder, providing robust semantic grounding and enabling effective perturbation-based augmentations. Empirically, we demonstrate that text representations can be compressed by 8×8$\times$8 × while maintaining generation quality comparable to token-level diffusion models. Furthermore, increasing the latent sequence length allows Cosmos to surpass both diffusion-based and autoregressive baselines. We evaluate Cosmos on four diverse generative tasks including story generation, question generation, summarization, and detoxification and compare it with various generative paradigms. Cosmos achieves comparable or superior generation quality while offering more than 2×2$\times$2 × faster inference.
@@ 1 Introduction
Motivated by these insights, we revisit the foundations of text representation. We hypothesize that the token-level encoding used in contemporary language models is heavily overparameterized for the purpose of sequence generation. Building on the successes of latent diffusion in vision, we pose the following question: How far can we compress textual information into a compact latent space while still matching or even exceeding the generative fidelity of conventional token-level representations?
@@
  • We propose Cosmos — a training recipe for a COmpressed and SMOoth latent Space, that allows to train a diffusion model in more compact latent space while matching the quality of token-level diffusion baselines.
  • We demonstrate that modestly scaling the number of latent vectors enables latent-space diffusion to surpass both token-level diffusion and autoregressive models on an unconditional text generation task.
  • On standard text-generation benchmarks, the proposed latent diffusion achieves up to $\mathbf{2\times}$ faster sampling than conventional token-level diffusion models, while matching or slightly exceeding them in quality and diversity.
@@ 4.1 Overview
We train a Gaussian diffusion model is space formed by $\mathbf{z}$, following Eqs. (1)–(2). Because $N$ is small, the diffusion model runs faster in this space.
@@ 4.2 Learning a compact text latent space
The previous Section 4.1 outlined our end-to-end pipeline: a frozen contextual encoder supplies token representations, a compressor distils them into a fixed-length latent matrix, and a diffusion model operates solely in this latent space. We now zoom in on the compression and decompression stages.
@@ 4.2.1 How compact can text latents be?
Figure 2: Token-level reconstruction accuracy on Wikipedia (512 tokens) as a function of $N$.
Figure 2 reports token-level reconstruction accuracy on Wikipedia dataset as a function of the latent sequence length $N$. The compressor outlined in §4.1 is trained using a token-level cross-entropy objective. Remarkably, a lightweight compressor equipped with only 12 transformer blocks is able to encode 512-token sequences into only $N=16$ latents with $100\%$ reconstruction accuracy, achieving a 32×32$\times$32 × compression relative to initial 512 hidden states produced by BERT.
These results paint an encouraging picture: if such a compact latent manifold can be modeled generatively, text generation could proceed in a space whose dimensionality is much smaller than that of the original token embeddings, promising dramatic speed-ups. However, as we show further, high reconstruction accuracy by itself does not imply that the latent manifold is suitable for generative modeling.
@@ 4.2.2 Learning robust compact text representations
The reconstruction study (§4.2.1) reveals that aggressive text compression can be loss-free. However, for a robust diffusion generation compressed space should satisfy additional requirements. Empirically, we observe that if the latent manifold lacks smoothness and robustness (Section 5), a Gaussian diffusion model fails to sample latents that yield high-quality texts. To improve diffusability of the latent space, we employ three complementary strategies to the autoencoder training.  Figure 1 provides an overview of the proposed approach.
@@
Alongside the cross-entropy loss between the original tokens and their reconstructions, we add a mean-squared error penalty between the frozen text encoder outputs $\mathbf{h}$ and their reconstructions $\mathbf{\hat{h}}$. This auxiliary objective forces the compressor to preserve the semantics carried by the contextual representations $\mathbf{h}$.
Activation-space perturbations.
@@
    Gaussian noise: after normalising $\mathbf{h}$ with the pre-computed statistics, we inject noise via $\mathbf{h}^{\prime}=\delta\,\mathbf{h}+\sqrt{1-\delta^{2}}\,\varepsilon$ with $\delta=0.7$ and $\varepsilon\sim\mathcal{N}(0,I)$.
These augmentations force the autoencoder to tolerate partial information loss and promote smooth interpolation between adjacent latents (see Section 5.1).
Latent-space augmentation.
We also apply augmentation directly to the latent matrix $\mathbf{z}$. Since the number of latent vectors $N$ can be an order of magnitude smaller than the token count $L$, masking whole latent vectors would annihilate too much information. Instead, during training we randomly zero out a fixed proportion $p$ of the individual features inside each latent vector. This fine-grained sparsification encourages neighbouring features within every latent vector to store redundant cues about one another, so that the representation remains intelligible even when a subset of features is excised, thereby making the manifold more robust to small latent perturbations..
5 Latent-space properties that facilitate diffusion training
This section embarks on an exploration of the text-latent manifold, asking which of its intrinsic features govern the performance of a diffusion model. All experiments reported here employ autoencoders that compress 128-token texts into 16 latent vectors. The autoencoders are trained on the Wikipedia dataset, whereas the diffusion models are optimised on the smaller yet high-quality ROCStories (mostafazadeh2016corpus, ) dataset. Across all experiments, we generate 1 000 samples and report each metric as the average over these samples. Our experiments spotlight two indispensable qualities — manifold smoothness and resilience to perturbations. The remainder of the section introduces an analysis that quantifies both attributes.
@@ 5.1 Smoothness of the latent manifold
    Apply diffusion noise (Equation 1) to get $\mathbf{z}^{\mu}_{t}$ at different time steps $t$ to measure how unseen regions influence diffusion at different noise levels.
@@
    Evaluate text plausibility with GPT–2 perplexity (PPL), averaged over 1 00010001\,0001 000 random endpoint pairs.
@@
By analysing how perplexity (PPL) evolves with the interpolation coefficient $\mu$ and the diffusion timestep $t$, we obtain a fine-grained picture of manifold smoothness and model robustness far beyond the training support. Figure 3 shows that autoencoder trained only with cross-entropy (CE) loss and our robustness-oriented alternative behave identically when $\mu\!\to\!0$ or $\mu\!\to\!1$, confirming that both models successfully reconstruct train-time input latents. As we move away from the endpoints, perplexity rises for both encoders, but it escalates far more rapidly for the CE baseline — clear evidence that the latent manifold learned by our autoencoder is markedly smoother.
@@ 5.2 Reducing the train–inference mismatch
To assess how well the decoder tolerates perturbations, we inject Gaussian noise into latent vectors of real texts: $\mathbf{z}_{\text{noised}}=\mathbf{z}+\sigma\,\boldsymbol{\varepsilon}$, where $\boldsymbol{\varepsilon}\sim\mathcal{N}(0,I)$, decode the perturbed vector to obtain $\mathbf{\hat{w}}$, and measure its BLEU score against the original text $\mathbf{w}$. Figure 4 shows that two components contribute most to decoder stability: (i) an explicit MSE loss on the decompressor’s output (§4.2.2), which prevents the final-layer norm from exploding, a common effect observed when training solely with cross-entropy loss (gao2021simcse, ), and (ii) latent masking (§4.2.2), which forces the compressor to distribute information more evenly among latent features and remain resilient to partial dropout. Notably, the decoder supplemented with described train-time modifications can reconstruct text almost perfectly even under substantial noise ($\sigma=1$).
@@
Next, we test the diffusion model’s sensitivity to small perturbations introduced mid-trajectory. For a late generation step $t$, chosen from the second half of the reverse process, where latent representations are already semantically meaningful, we add Gaussian noise with magnitude $\nu$ and continue sampling. The deviation from the original trajectory is quantified by the MSE between the final estimates, $\lVert\hat{\mathbf{z}}-\hat{\mathbf{z}}^{\text{shifted}}\rVert_{2}^{2}$.
@@
Configuration MSE$\bigl{(}\hat{\mathbf{z}},E(D(\hat{\mathbf{z}}))\bigr{)}$
$\mathrm{CE}(w,\hat{w})$ 0.721
@@
Table 1 tracks the train–inference mismatch, quantified as the MSE between $\hat{\mathbf{z}}$ and $E\bigl{(}D(\hat{\mathbf{z}})\bigr{)}$. Each proposed refinement tightens the gap. The consistent drop in error confirms that the enhanced training procedure makes the autoencoder more resilient to diffusion perturbations. Together, these results demonstrate that our training recipe simultaneously strengthens the decoder and the diffusion model, thereby shrinking the train–inference gap.
@@ 6.1 Empirical analysis of autoencoder training regimes
We evaluate the unconditional generation capacity of models with three complementary metrics that, together, capture textual quality, lexical diversity, and distributional fidelity. First, we report perplexity (PPL), computed with GPT-2 Large (radford2019language, ). Second, we quantify diversity using the score $\mathrm{\textbf{div}}(y)=\prod_{n=2}^{4}\tfrac{\#\,\text{unique }n\text{-gramsin }y}{\#\,n\text{-grams in }y}$, where $y$ denotes the set of generated texts. Finally, because low perplexity can be achieved by simple or repetitive texts, we also compute the MAUVE score (pillutla2021mauve, ), which measures the distributional alignment between generated and reference texts and thus offers a broader view of generation quality. For every model we generate $1{,}000$ samples and, when calculating MAUVE, draw an equally sized reference subset from the held-out test split. The entire evaluation procedure is repeated ten times with different random seeds, and we report the mean and standard deviation of the resulting scores across runs.
@@ Results.
Table 2 demonstrates that each proposed feature contributes significantly to the model performance. Remarkably, adding MSE penalty between the predicted and reference BERT activations alone boosts all quality metrics by about 50%. Subsequent augmentation of the BERT activations yields a further leap, underscoring the value of intermediate contextual, semantically grounded representations. This stage rises MAUVE to 0.767, already achieving the performance obtained when the diffusion model is trained on the full-length, uncompressed BERT representations (0.767 vs. 0.762 for TEncDM (shabalin2025tencdm, )). Finally, dropping random features from latent vectors pushes MAUVE beyond 0.83, lowers perplexity to 30.2, keeping diversity essentially unchanged.
These results provide a clear answer to the question posed in the introduction of this study: text representations can indeed be mapped into a more compact latent space, where a trained diffusion model performs comparably to traditional token-level counterparts without sacrificing the text generation quality. For all subsequent autoencoders we lock in the configuration highlighted in Table 2: the MSE penalty is kept, the random-masking rate is fixed at $0.3$, Gaussian noising is applied with $\delta=0.7$, and latent masking rate is set to $0.4$.
Additionally, we validate all proposed autoencoder configurations by training the diffusion model on the larger Wikipedia dataset. Detailed results of this evaluation are presented in Section C.2. Furthermore, we explore the use of a variational prior in the latent space and observe that it offers no clear advantage. A thorough analysis is provided in Section C.1.
@@ 6.2 Impact of scaling the number of latent vectors
Figure 6: Impact of scaling $N$ on diffusion generation quality.
In this section, we examine how diffusion generation quality varies when varying the number of latent vectors $N$. We keep the training pipeline fixed according to the hyperparameters detailed in Section 6.1. We do not modify the embedding dimension $d$, because it would necessitate architectural modifications to the diffusion model, making it difficult to isolate the impact of latent space diffusability from changes in model capacity and structure.
Table 3: Impact of scaling the number of latent vectors $N$ on unconditional generation quality.
$N$ MAUVE $\uparrow$ PPL $\downarrow$ Div $\uparrow$
@@
Figure 6 compares baseline cross-entropy compressor to the robustness-oriented compressor introduced in Section 4.2.2. We observe that for CE compressor decrease of $N$ leads to a rapid degradation in quality. In contrast, the proposed autoencoder maintains high generation quality even under substantial compression. It achieves an 8×8$\times$8 × reduction in latent sequence length while surpassing the quality of uncompressed representations.
Table 8 provides further insights by comparing generation quality across different numbers of latent vectors $N$. Although configuration $N=16$ matches the quality of BERT representations baseline, we observe a notable quality gap compared to the configuration with $N{=}128$: the MAUVE score decreases from $0.940$ to $0.836$, while perplexity increases from $26.3$ to $30.2$. However, adopting a less aggressive compression to $N{=}32$ (a 4×4$\times$4 × compression) results in only minor quality degradation, still significantly outperforming the baseline. Based on these observations, together with results from Section 7, we recommend limiting latent compression to approximately 4×4$\times$4 × to optimally balance quality and computational efficiency.
@@ 7 Comparison across generative paradigms
We evaluate the performance of Cosmos on four distinct text generation tasks: unconditional story generation (ROCStories (mostafazadeh2016corpus, )), text summarization (XSum (narayan2018don, )), detoxification (ParaDetox (logacheva-etal-2022-paradetox, )), and question generation (SQuAD2.0 (minaee2021deep, )). For unconditional story generation we employ metrics discussed in Section 6.1. For text summarization and question generation, we report BERTScore (BS) (zhang2019bertscore, ). The detoxification task is evaluated using BLEU-4 (BLEU). We further assess models’ efficiency by comparing inference time for unconditional generation on sequences with 128 and 512 tokens. Additional metrics and implementation details are provided in the Section C.3.
@@
[…]
@@
Method MAUVE $\uparrow$ PPL $\downarrow$ Div $\uparrow$ BS $\uparrow$ BLEU $\uparrow$ BS $\uparrow$ $L=128$ $L=512$
Source text 0.953 21.7 0.403
GPT2 0.789 20.5 0.252 0.690 0.677 0.680 $\mathbf{1.2_{.1}}$ $42.8_{.1}$
GPT Neo 0.720 19.9 0.258 0.621 0.610 0.665 $2.6_{.1}$ $45.2_{.2}$
AR-Diffusion 0.066 41.8 0.101 0.568 0.647 0.569 $226.4_{.2}$ $602_{2.1}$
DiffuSeq 0.086 50.5 0.124 0.588 0.679 0.563 $215.9_{1.2}$ $1565_{.7}$
SeqDiffuSeq 0.103 29.3 0.137 0.617 0.688 0.574 $100_{.5}$ $601_{.5}$
TESS 0.061 22.4 0.163 0.627 0.693 0.667 $1441_{.2}$ $5984_{.0}$
SEDD 0.598 70.8 0.336 0.576 0.666 0.443 $15.0_{1.1}$ $60.3_{1.0}$
LD4LG 0.716 30.6 0.331 0.702 0.708 0.641 $27.9_{.5}$ $102_{.2}$
TEncDM 0.762 29.1 0.295 0.699 0.619 0.703 $29.6_{.1}$ $180_{.3}$
CosmosN=16 0.836 30.2 0.322 0.654 $5.8_{.1}$
CosmosN=128 0.940 26.3 0.346 0.704 0.694 0.708 $35.1_{.1}$ $\mathbf{36.6_{.1}}$

### v2

@@ Abstract
Autoregressive language models dominate modern text generation, yet their sequential nature introduces fundamental limitations: decoding is slow, and maintaining global coherence remains challenging. Diffusion models offer a promising alternative by enabling parallel generation and flexible control; however, their application to text generation is hindered by the high dimensionality of token-level representations. We introduce Cosmos, a novel approach to text generation that operates entirely in a compressed, smooth latent space tailored specifically for diffusion. This space is learned using an autoencoder trained simultaneously for token-level reconstruction and alignment with frozen activations from a pretrained language encoder, providing robust semantic grounding and enabling effective perturbation‑based augmentations. Empirically, we demonstrate that text representations can be compressed up to $8\times$ while maintaining generation quality comparable to token‑level diffusion models. Furthermore, increasing the latent sequence length allows Cosmos to surpass both diffusion‑based and autoregressive baselines. We evaluate Cosmos on four diverse generative tasks including story generation, question generation, summarization, and detoxification and compare it with various generative paradigms. Cosmos achieves comparable or superior generation quality while offering more than $2\times$ faster inference. Code is released at GitHub.
11footnotetext: Corresponding author: vmeshchaninov@hse.ru
@@ 1 Introduction
Motivated by these insights, we revisit the foundations of text representation. We hypothesize that the token-level encoding used in contemporary language models is heavily overparameterized for the purpose of sequence generation. Building on the successes of latent diffusion in vision, we pose the following question: How far can we compress textual information into a compact latent space while still matching or even exceeding the generative fidelity of conventional token‑level representations?
@@
  • We propose Cosmos — a training recipe for a COmpressed and SMOoth latent Space, that allows to train a diffusion model in more compact latent space while matching the quality of token‑level diffusion baselines.
  • We demonstrate that modestly scaling the number of latent vectors enables latent‑space diffusion to surpass both token‑level diffusion and autoregressive models on an unconditional text generation task.
  • On standard text‑generation benchmarks, the proposed latent diffusion achieves up to $\mathbf{2\times}$ faster sampling than conventional token‑level diffusion models, while matching or slightly exceeding them in quality and diversity.
@@ 2 Related work
Our study builds upon the approaches reviewed here. For additional context and a discussion of related topics please see Appendix B.
@@ 4.1 Overview
We train a Gaussian diffusion model is space formed by $\mathbf{z}$, following Eqs. (1)–(2). Because NN is small, the diffusion model runs faster in this space.
@@ 4.2 Learning a compact text latent space
The previous Section˜4.1 outlined our end‑to‑end pipeline: a frozen contextual encoder supplies token representations, a compressor distils them into a fixed‑length latent matrix, and a diffusion model operates solely in this latent space. We now zoom in on the compression and decompression stages.
@@ 4.2.1 How compact can text latents be?
Figure 2: Token‑level reconstruction accuracy on Wikipedia (512 tokens) as a function of $N$.
Figure˜2 reports token‑level reconstruction accuracy on Wikipedia dataset as a function of the latent sequence length $N$. The compressor outlined in §4.1 is trained using a token‑level cross‑entropy objective. Remarkably, a lightweight compressor equipped with only 12 transformer blocks is able to encode 512‑token sequences into only $N=16$ latents with $100\%$ reconstruction accuracy, achieving a $32\times$ compression relative to initial 512 hidden states produced by BERT.
These results paint an encouraging picture: if such a compact latent manifold can be modeled generatively, text generation could proceed in a space whose dimensionality is much smaller than that of the original token embeddings, promising dramatic speed‑ups. However, as we show further, high reconstruction accuracy by itself does not imply that the latent manifold is suitable for generative modeling.
@@ 4.2.2 Learning robust compact text representations
The reconstruction study (§4.2.1) reveals that aggressive text compression can be loss‑free. However, for a robust diffusion generation compressed space should satisfy additional requirements. Empirically, we observe that if the latent manifold lacks smoothness and robustness (Section˜5), a Gaussian diffusion model fails to sample latents that yield high‑quality texts. To improve diffusability of the latent space, we employ three complementary strategies to the autoencoder training.  Figure˜1 provides an overview of the proposed approach.
@@
Alongside the cross‑entropy loss between the original tokens and their reconstructions, we add a mean‑squared error penalty between the frozen text encoder outputs $\mathbf{h}$ and their reconstructions $\mathbf{\hat{h}}$. This auxiliary objective forces the compressor to preserve the semantics carried by the contextual representations $\mathbf{h}$.
Activation‑space perturbations.
@@
    Gaussian noise: after normalising $\mathbf{h}$ with the pre‑computed statistics, we inject noise via $\mathbf{h}^{\prime}=\delta\,\mathbf{h}+\sqrt{1-\delta^{2}}\,\varepsilon$ with $\delta=0.7$ and $\varepsilon\sim\mathcal{N}(0,I)$.
These augmentations force the autoencoder to tolerate partial information loss and promote smooth interpolation between adjacent latents (see Section˜5.1).
Latent‑space augmentation.
We also apply augmentation directly to the latent matrix $\mathbf{z}$. Since the number of latent vectors $N$ can be an order of magnitude smaller than the token count $L$, masking whole latent vectors would annihilate too much information. Instead, during training we randomly zero out a fixed proportion $p$ of the individual features inside each latent vector. This fine‑grained sparsification encourages neighbouring features within every latent vector to store redundant cues about one another, so that the representation remains intelligible even when a subset of features is excised, thereby making the manifold more robust to small latent perturbations..
5 Latent‑space properties that facilitate diffusion training
This section embarks on an exploration of the text‑latent manifold, asking which of its intrinsic features govern the performance of a diffusion model. All experiments reported here employ autoencoders that compress 128‑token texts into 16 latent vectors. The autoencoders are trained on the Wikipedia dataset, whereas the diffusion models are optimised on the smaller yet high-quality ROCStories (mostafazadeh2016corpus, ) dataset. Across all experiments, we generate 1 000 samples and report each metric as the average over these samples. Our experiments spotlight two indispensable qualities — manifold smoothness and resilience to perturbations. The remainder of the section introduces an analysis that quantifies both attributes.
@@ 5.1 Smoothness of the latent manifold
    Apply diffusion noise (Equation˜1) to get $\mathbf{z}^{\mu}_{t}$ at different time steps $t$ to measure how unseen regions influence diffusion at different noise levels.
@@
    Evaluate text plausibility with GPT–2 perplexity (PPL), averaged over 1 0001\,000 random endpoint pairs.
@@
By analysing how perplexity (PPL) evolves with the interpolation coefficient $\mu$ and the diffusion timestep $t$, we obtain a fine‑grained picture of manifold smoothness and model robustness far beyond the training support. Figure 3 shows that autoencoder trained only with cross‑entropy (CE) loss and our robustness‑oriented alternative behave identically when $\mu\!\to\!0$ or $\mu\!\to\!1$, confirming that both models successfully reconstruct train-time input latents. As we move away from the endpoints, perplexity rises for both encoders, but it escalates far more rapidly for the CE baseline — clear evidence that the latent manifold learned by our autoencoder is markedly smoother.
@@ 5.2 Reducing the train–inference mismatch
To assess how well the decoder tolerates perturbations, we inject Gaussian noise into latent vectors of real texts: $\mathbf{z}_{\text{noised}}=\mathbf{z}+\sigma\,\boldsymbol{\varepsilon}$, where $\boldsymbol{\varepsilon}\sim\mathcal{N}(0,I)$, decode the perturbed vector to obtain $\mathbf{\hat{w}}$, and measure its BLEU score against the original text $\mathbf{w}$. Figure 4 shows that two components contribute most to decoder stability: (i) an explicit MSE loss on the decompressor’s output (§4.2.2), which prevents the final‑layer norm from exploding, a common effect observed when training solely with cross-entropy loss (gao2021simcse, ), and (ii) latent masking (§4.2.2), which forces the compressor to distribute information more evenly among latent features and remain resilient to partial dropout. Notably, the decoder supplemented with described train-time modifications can reconstruct text almost perfectly even under substantial noise ($\sigma=1$).
@@
Next, we test the diffusion model’s sensitivity to small perturbations introduced mid‑trajectory. For a late generation step $t$, chosen from the second half of the reverse process, where latent representations are already semantically meaningful, we add Gaussian noise with magnitude $\nu$ and continue sampling. The deviation from the original trajectory is quantified by the MSE between the final estimates, $\lVert\hat{\mathbf{z}}-\hat{\mathbf{z}}^{\text{shifted}}\rVert_{2}^{2}$.
@@
[…]
@@ 6.1 Empirical analysis of autoencoder training regimes
[…]
@@ Results.
Table˜2 demonstrates that each proposed feature contributes significantly to the model performance. Remarkably, adding MSE penalty between the predicted and reference BERT activations alone boosts all quality metrics by about 50%. Subsequent augmentation of the BERT activations yields a further leap, underscoring the value of intermediate contextual, semantically grounded representations. This stage rises MAUVE to 0.767, already achieving the performance obtained when the diffusion model is trained on the full‑length, uncompressed BERT representations (0.767 vs. 0.762 for TEncDM (shabalin2025tencdm, )). Finally, dropping random features from latent vectors pushes MAUVE beyond 0.83, lowers perplexity to 30.2, keeping diversity essentially unchanged.
These results provide a clear answer to the question posed in the introduction of this study: text representations can indeed be mapped into a more compact latent space, where a trained diffusion model performs comparably to traditional token-level counterparts without sacrificing the text generation quality. For all subsequent autoencoders we lock in the configuration highlighted in Table˜2: the MSE penalty is kept, the random‐masking rate is fixed at $0.3$, Gaussian noising is applied with $\delta=0.7$, and latent masking rate is set to $0.4$.
Additionally, we validate all proposed autoencoder configurations by training the diffusion model on the larger Wikipedia dataset. Detailed results of this evaluation are presented in LABEL:app:ablation_wikipedia. Furthermore, we explore the use of a variational prior in the latent space and observe that it offers no clear advantage. A thorough analysis is provided in Section˜D.1.
@@ 6.2 Impact of scaling the number of latent vectors
Figure 6: Impact of scaling NN on diffusion generation quality.
In this section, we examine how diffusion generation quality varies when varying the number of latent vectors $N$. We keep the training pipeline fixed according to the hyperparameters detailed in Section˜6.1. We do not modify the embedding dimension $d$, because it would necessitate architectural modifications to the diffusion model, making it difficult to isolate the impact of latent space diffusability from changes in model capacity and structure.
Table 3: Impact of scaling the number of latent vectors NN on unconditional generation quality.
NN MAUVE $\uparrow$ PPL $\downarrow$ Div $\uparrow$
@@
Figure˜6 compares baseline cross-entropy compressor to the robustness-oriented compressor introduced in Section˜4.2.2. We observe that for CE compressor decrease of $N$ leads to a rapid degradation in quality. In contrast, the proposed autoencoder maintains high generation quality even under substantial compression. It achieves an $8\times$ reduction in latent sequence length while surpassing the quality of uncompressed representations.
Table 3 provides further insights by comparing generation quality across different numbers of latent vectors $N$. Although configuration $N=16$ matches the quality of BERT representations baseline, we observe a notable quality gap compared to the configuration with N=128N{=}128: the MAUVE score decreases from $0.940$ to $0.836$, while perplexity increases from $26.3$ to $30.2$. However, adopting a less aggressive compression to N=32N{=}32 (a $4\times$ compression) results in only minor quality degradation, still significantly outperforming the baseline. Based on these observations, together with results from Section˜7, we recommend limiting latent compression to approximately $4\times$ to optimally balance quality and computational efficiency.
@@ 7 Comparison across generative paradigms
We evaluate the performance of Cosmos on four distinct text generation tasks: unconditional story generation (ROCStories (mostafazadeh2016corpus, )), text summarization (XSum (narayan2018don, )), detoxification (ParaDetox (logacheva-etal-2022-paradetox, )), and question generation (SQuAD2.0 (minaee2021deep, )). For unconditional story generation we employ metrics discussed in Section˜6.1. For text summarization and question generation, we report BERTScore (BS) (zhang2019bertscore, ). The detoxification task is evaluated using BLEU-4 (BLEU). We further assess models’ efficiency by comparing inference time for unconditional generation on sequences with 128 and 512 tokens. Additional metrics and implementation details are provided in the Section˜D.2.
@@
To perform latent-space diffusion, we begin by training autoencoders on Wikipedia dataset, varying the input sequence length. For tasks with relatively short input, such as unconditional story generation and text detoxification, we design two variants of the autoencoder that map 128-token input into either $16$ or $128$ latent vectors, resulting in the CosmosN=16N=16 and CosmosN=128N=128 configurations, respectively. In contrast, conditional generation tasks such as summarization and question generation require substantially longer input contexts. For these settings, we employ a third autoencoder configured to compress 512-token sequences into 128 latent vectors. Accordingly, we do not report results for long-context tasks for CosmosN=16N=16, as sych aggressive compression fails to yield plausible outputs in these long-context scenarios. We adopt N=128N{=}128 ($4\times$ compression) for long-context tasks as it strikes a favorable balance between generation quality and computational efficiency, as demonstrated in the scaling analysis in Section˜6.2.
@@
Method MAUVE $\uparrow$ PPL $\downarrow$ Div $\uparrow$ Mem $\downarrow$ BS $\uparrow$ BLEU $\uparrow$ BS $\uparrow$ $L=128$ $L=512$
Source text 0.953 21.7 0.403 0.365
GPT2 0.789 20.5 0.252 0.455 0.690 0.677 0.680 $\mathbf{1.2_{.1}}$ $42.8_{.1}$
GPT Neo 0.720 19.9 0.258 0.469 0.621 0.610 0.665 $2.6_{.1}$ $45.2_{.2}$
AR-Diffusion 0.066 41.8 0.101 0.540 0.568 0.647 0.569 $226.4_{.2}$ $602_{2.1}$
DiffuSeq 0.086 50.5 0.124 0.516 0.588 0.679 0.563 $215.9_{1.2}$ $1565_{.7}$
SeqDiffuSeq 0.103 29.3 0.137 0.663 0.617 0.688 0.574 $100_{.5}$ $601_{.5}$
TESS 0.061 22.4 0.163 0.550 0.627 0.693 0.667 $1441_{.2}$ $5984_{.0}$
SEDD 0.598 70.8 0.336 0.325 0.576 0.666 0.443 $15.0_{1.1}$ $60.3_{1.0}$
LD4LG 0.716 30.6 0.331 0.432 0.702 0.708 0.641 $27.9_{.5}$ $102_{.2}$
TEncDM 0.762 29.1 0.295 0.438 0.699 0.619 0.703 $29.6_{.1}$ $180_{.3}$
CosmosN=16 0.836 30.2 0.322 0.394 0.654 $5.8_{.1}$
CosmosN=128 0.940 26.3 0.346 0.383 0.704 0.694 0.708 $35.1_{.1}$ $\mathbf{36.6_{.1}}$